# Levi-Civita connection on the irreducible quantum flag manifolds

### v1

@@ 1. Introduction
The article is organised as follows: in Section 2, we recall some necessary preliminaries about differential calculi, metrics (in the sense of Beggs and Majid), connections and their compatibility with metrics, bar categories and noncommutative complex structures.
@@ 2. Preliminaries
Let $\operatorname{ev}_{V}:{}^{\ast}V\otimes V\to\mathbb{C}$ be a choice of evaluation map for $V$, and denote the uniquely determined right coevaluation map by $\operatorname{coev}_{V}$. Moreover, choose an evaluation map $\operatorname{ev}_{{}^{\ast}V}:V\otimes{}^{\ast}V\to\mathbb{C}$, and denote the uniquely determined coevaluation map by $\operatorname{coev}_{{}^{\ast}V}$. Then a self-dual structure for $V\oplus{}^{\ast}V$ is given by
$\left(\operatorname{ev}:=\operatorname{ev}_{V}\oplus\operatorname{ev}_{{}^{\ast}V},\,\operatorname{coev}:=\operatorname{coev}_{V}\oplus\operatorname{coev}_{{}^{\ast}V}\right)$
@@ 2.1. Hopf-algebras and relative Hopf-modules
[Ulb, Lemma 2.2] implies that $(\prescript{}{B}{\operatorname{Hom}}(M,B),\prescript{\prescript{}{B}{\operatorname{Hom}}(M,B)}{}{\delta})$ is a comodule while [OBvvR, Proposition A.9] proves that $\prescript{}{B}{\operatorname{Hom}}(M,B)$ is an object in ${{}^{\phantom{B}\mathllap{A}}_{\phantom{B}\mathllap{B}}\mathsf{Mod}_{B}}$. It is well known that the maps $\operatorname{ev},\operatorname{coev}$ defined in (5) are morphisms in ${{}^{\phantom{B}\mathllap{}}_{\phantom{B}\mathllap{B}}\mathsf{Mod}_{B}}$. By a routine computation, it can be checked that $\operatorname{ev}$ and $\operatorname{coev}$ are also left $A$-colinear. ∎
@@ 2.2. Takeuchi’s equivalence
$M/{B^{+}M}\rightarrow H\otimes M/{B^{+}M},\leavevmode\nobreak\ [m]\mapsto\pi(m_{(-1)})\otimes[m_{(0)}]$
is a coaction of $H$ on $M/{B^{+}M}$, where $[m]$ denotes the equivalence class of $m\in M$ in $M/{B^{+}M}$. Thus, $M/{B^{+}M}$ is an object of the category ${{}^{\phantom{H}\mathllap{H}}_{\phantom{H}\mathllap{}}\mathsf{Mod}_{B}}$ with respect to the obvious right $B$-module structure and we have a functor
$\Phi:{{}^{\phantom{B}\mathllap{A}}_{\phantom{B}\mathllap{B}}\mathsf{Mod}_{B}}\rightarrow{{}^{\phantom{H}\mathllap{H}}_{\phantom{H}\mathllap{}}\mathsf{Mod}_{B}},\leavevmode\nobreak\ M\rightarrow M/B^{+}M.$ (9)
@@
The categories ${}^{\phantom{B}\mathllap{A}}_{\phantom{B}\mathllap{B}}\mathsf{mod}_{0}$ and ${{}^{\phantom{H}\mathllap{H}}_{\phantom{H}\mathllap{}}\mathsf{mod}}$ are monoidal, the monoidal structure on ${}^{\phantom{B}\mathllap{A}}_{\phantom{B}\mathllap{B}}\mathsf{mod}_{0}$ is the usual tendor product of comodules and $B$-bimodules while the monoidal structure on ${{}^{\phantom{H}\mathllap{H}}_{\phantom{H}\mathllap{}}\mathsf{mod}}$ is given by the usual tensor product of comodules.
@@ 2.3. Differential calculi and metrics
If $(g,(\leavevmode\nobreak\ ,\leavevmode\nobreak\ ))$ is a metric on $B$, then by [BM4, Lemma 1.16], $g$ is central, i.e, $bg=gb$ for all $b\in B$. This implies that the map $\operatorname{coev}_{g}$ is actually $B$-bilinear. The following well-known characterisation (see page 311 of [BM4]) of metrics will be used throughout the article repeatedly.
@@
If $(\Omega^{\bullet},\wedge,d)$ is a *-differential calculus on *-algbera $B$ and $(g,(\leavevmode\nobreak\ ,\leavevmode\nobreak\ ))$ is metric on $\Omega^{1}$ then $g^{\dagger}=g$ if and only if $(\omega,\eta)=(\eta^{*},\omega^{*})^{*}$ for all $\omega,\eta\in\Omega^{1}$.
@@ 2.4. Complex structures
Following the work of Polishchuk and Schwarz in [PS], the notion of complex structures for noncommutative differential calculus have been studied by many mathematicians, see for example, [KLvS], [BS], and references therein. We will work with the following definition introduced in [OB1].
@@ 2.5. Bar Categories
A bar category is a monoidal category $(\mathcal{C},\otimes,1_{\mathcal{C}})$ together with a functor $\mathrm{bar}:\mathcal{C}\to\mathcal{C}$ (written as $X\mapsto\overline{X}$), a natural equivalence $\mathrm{bb}:\mathrm{id}_{\mathcal{C}}\to\mathrm{bar}\circ\mathrm{bar}$ between the identity and the bar $\circ$ bar functors on $\mathcal{C}$, an invertible morphism $\star:1_{\mathcal{C}}\to\overline{1_{\mathcal{C}}}$ and a natural equivalence $\Upsilon$ (capital upsilon) between $\mathrm{bar}\circ\otimes$ and $\otimes\circ(\mathrm{bar}\times\mathrm{bar})\circ\mathrm{flip}$ from $\mathcal{C}\times\mathcal{C}$ to $\mathcal{C}$
@@
and moreover, the following equations hold
@@
An object $X$ in a bar category is called a star object if there is morphism $\star_{X}:X\to\overline{X}$ such that $\overline{\star_{X}}\circ\star_{X}=\mathrm{bb}_{X}$.
@@ 2.6. Connections
([BM4]) Suppose $g$ is a metric on the space of one-forms $\Omega^{1}$ of a differential calculus on an algebra $B$ in the sense of Definition 2.9. Then a bimodule connection $(\nabla,\sigma)$ is said to be compatible with $g$ if $\nabla g=0$, where, by an abuse of notation, we have denoted the connection on $\Omega^{1}\otimes_{B}\Omega^{1}$ defined by (16) by $\nabla$ again.
@@ 3. Construction of metrics on quantum homogeneous spaces
In this section, we start by classifying covariant metrics on a quantum homogeneous space of the form $B=A^{\mathrm{co}(H)},$ where $A$ is a Hopf $\ast$-algebra and $H$ is a compact quantum group algebra (see Subsection 8.1). This is proved in Theorem 3.2. We will use this result to classify covariant metrics on the space of one-forms of the Heckenberger–Kolb calculus in Theorem 6.12. In the second subsection, we work in the set up of covariant complex structures on a quantum homogeneous space as above. Here, we construct a family of covariant metrics $(g,(\leavevmode\nobreak\ ,\leavevmode\nobreak\ ))$ which are real in the sense of Definition 2.12 and satisfies the condition $(\omega,\eta)=0$ if $\omega,\eta\in\Omega^{(1,0)}$ or if $\omega,\eta\in\Omega^{(0,1)}$.
@@ 3.1. The moduli space of metrics
So let us assume that $H$ is a compact quantum group algebra and moreover, $\Omega^{1}$ be the space of one-forms of a covariant $\ast$-differential calculus on $B$ such that $\Omega^{1}$ is an object of the category ${}^{\phantom{B}\mathllap{A}}_{\phantom{B}\mathllap{B}}\mathsf{mod}_{0}$. We will use the notation $V$ to denote $\Phi(\Omega^{1})$. Since the map $\star_{\Omega^{1}}$ is an isomorphism in ${}^{\phantom{B}\mathllap{A}}_{\phantom{B}\mathllap{B}}\mathsf{mod}_{0}$, we know that $\Phi(\star_{\Omega^{1}})$ induces an isomorphism in ${{}^{\phantom{H}\mathllap{H}}_{\phantom{H}\mathllap{}}\mathsf{mod}}$ from $V$ to $\overline{V}$ via Lemma 8.7. Moreover, if we fix an $H$-invariant inner product on $V$ (conjugate-linear on the right) in the sense of (57), then we have the isomorphism $\psi:\overline{V}\to\operatorname{Hom}(V,\mathbb{C})$ in ${{}^{\phantom{H}\mathllap{H}}_{\phantom{H}\mathllap{}}\mathsf{mod}}$ from Lemma 8.1. Hence, we have a right dual pairing of $V$ with itself defined by the following morphisms in ${{}^{\phantom{H}\mathllap{H}}_{\phantom{H}\mathllap{}}\mathsf{mod}}$:
@@
Now, by Takeuchi’s equivalence (Theorem 2.6), we have morphisms $\operatorname{ev}:\Omega^{1}\otimes_{B}\Omega^{1}\to B$ and $\operatorname{coev}:B\to\Omega^{1}\otimes_{B}\Omega^{1}$ in ${}^{\phantom{B}\mathllap{A}}_{\phantom{B}\mathllap{B}}\mathsf{mod}_{0}$ where $\Phi(\operatorname{ev})$ and $\Phi(\operatorname{coev})$ are defined by (17) and (18) respectively. Thus, we immediately have the following result.
@@
Since $H$ is a compact quantum group algebra, the set of all $A$-covariant metrics on $\Omega^{1}$ is non-empty by Proposition 3.1. Consider the monoidal category ${}^{\phantom{B}\mathllap{A}}_{\phantom{B}\mathllap{B}}\mathsf{mod}_{0}$. Then by a combination of Remark 2.10 and Lemma 2.1, it follows that the space of $A$-covariant metrics is acted upon freely and transitively by the automorphism group of $\Omega^{1}$ in the category ${}^{\phantom{B}\mathllap{A}}_{\phantom{B}\mathllap{B}}\mathsf{mod}_{0}$. However, by Takuechi’s equivalence, this group is isomorphic to the automorphism group of $\Phi(\Omega^{1})$ in the category ${{}^{\phantom{H}\mathllap{H}}_{\phantom{H}\mathllap{}}\mathsf{mod}}$.
@@ 3.2. Metrics on quantum homogeneous spaces with a complex structure
Using the fact that $MB^{+}=B^{+}M$ for any object $M$ in ${}^{\phantom{B}\mathllap{A}}_{\phantom{B}\mathllap{B}}\mathsf{mod}_{0}$ (see Definition 2.5), it can be easily checked that the map $\operatorname{ev}_{V^{(1,0)}}$ is well-defined.
@@
as $(\leavevmode\nobreak\ ,\leavevmode\nobreak\ )_{\Omega^{1}}=(\leavevmode\nobreak\ ,\leavevmode\nobreak\ )^{1,0}\oplus(-\lambda^{-1})(\leavevmode\nobreak\ ,\leavevmode\nobreak\ )^{0,1}\oplus 0\oplus 0$ and $\operatorname{coev}_{\Omega^{1}}=\operatorname{coev}^{1,0}\oplus(-\lambda)\operatorname{coev}^{0,1}\oplus 0\oplus 0$, where $(\leavevmode\nobreak\ ,\leavevmode\nobreak\ )^{1,0}$, $\operatorname{coev}^{1,0}$ and $(\leavevmode\nobreak\ ,\leavevmode\nobreak\ )^{0,1},\operatorname{coev}^{0,1}$ are as defined in Corollary 3.4 and Remark 3.5 respectively. If we denote $\operatorname{coev}_{\Omega^{1}}(1)$ by the symbol $g$, then $(g,(\leavevmode\nobreak\ ,\leavevmode\nobreak\ )_{\Omega^{1}})$ defines an $A$-covariant real metric on $\Omega^{1}$ (see Definition 2.12) which satisfies the following conditions
@@ 4.1. Construction of Hermitian metrics on quantum homogeneous spaces
Since the element $\sum^{n}_{i=1}e_{i}\otimes_{B}e^{i}$ is the inverse image of $\mathrm{id}_{\Omega^{1}}$ under the canonical isomorphism $\prescript{}{B}{\operatorname{Hom}(\Omega^{1},B)}\rightarrow\prescript{}{B}{\operatorname{Hom}(\Omega^{1},\Omega^{1})},$ the definition of $g$ is independent of the choice of the dual basis. It can be easily checked that $(g,(\leavevmode\nobreak\ ,\leavevmode\nobreak\ ))$ is a covariant metric on $\Omega^{1}.$
@@
Suppose $(\Omega^{(\bullet,\bullet)},\wedge,\partial,\overline{\partial})$ be an $A$-covariant complex structure on a quantum homogeneous space $B=A^{\mathrm{co}(H)}$, where $A$ is a Hopf $\ast$-algebra and $\Omega^{1}$ is an object of ${}^{\phantom{B}\mathllap{A}}_{\phantom{B}\mathllap{B}}\mathsf{mod}_{0}$. Moreover, we assume that $\Omega^{(1,0)}$ and $\Omega^{(0,1)}$ are non isomorphic simple objects in ${}^{\phantom{B}\mathllap{A}}_{\phantom{B}\mathllap{B}}\mathsf{mod}_{0}$.

### v2

@@ 1. Introduction
The article is organized as follows: in Section 2, we recall some necessary preliminaries about differential calculi, metrics (in the sense of Beggs and Majid), connections and their compatibility with metrics, bar categories and noncommutative complex structures.
@@ 2. Preliminaries
Let $\operatorname{ev}_{V}:V\otimes{}^{\ast}V\to\mathbb{C}$ be a choice of evaluation map for $V$, and denote the uniquely determined right coevaluation map by $\operatorname{coev}_{V}$. Moreover, choose an evaluation map $\operatorname{ev}_{{}^{\ast}V}:{}^{\ast}V\otimes V\to\mathbb{C}$, and denote the uniquely determined coevaluation map by $\operatorname{coev}_{{}^{\ast}V}$. Then a self-dual structure for $V\oplus{}^{\ast}V$ is given by
[…]
@@ 2.1. Hopf-algebras and relative Hopf-modules
[Ulb, Lemma 2.2] implies that $(\prescript{}{B}{\operatorname{Hom}}(M,B),\prescript{\prescript{}{B}{\operatorname{Hom}}(M,B)}{}{\delta})$ is a comodule while [OSvR, Proposition A.9] proves that $\prescript{}{B}{\operatorname{Hom}}(M,B)$ is an object in ${{}^{\phantom{B}\mathllap{A}}_{\phantom{B}\mathllap{B}}\mathsf{Mod}_{B}}$. It is well known that the maps $\operatorname{ev},\operatorname{coev}$ defined in (5) are morphisms in ${{}^{\phantom{B}\mathllap{}}_{\phantom{B}\mathllap{B}}\mathsf{Mod}_{B}}$. By a routine computation, it can be checked that $\operatorname{ev}$ and $\operatorname{coev}$ are also left $A$-colinear. ∎
@@ 2.2. Takeuchi’s equivalence
$\frac{M}{B^{+}M}\rightarrow H\otimes\frac{M}{B^{+}M},\leavevmode\nobreak\ [m]\mapsto\pi(m_{(-1)})\otimes[m_{(0)}]$
is a coaction of $H$ on $\frac{M}{B^{+}M}$, where $[m]$ denotes the equivalence class of $m\in M$ in $\frac{M}{B^{+}M}$. Thus, $\frac{M}{B^{+}M}$ is an object of the category ${{}^{\phantom{H}\mathllap{H}}_{\phantom{H}\mathllap{}}\mathsf{Mod}_{B}}$ with respect to the obvious right $B$-module structure and we have a functor
$\Phi:{{}^{\phantom{B}\mathllap{A}}_{\phantom{B}\mathllap{B}}\mathsf{Mod}_{B}}\rightarrow{{}^{\phantom{H}\mathllap{H}}_{\phantom{H}\mathllap{}}\mathsf{Mod}_{B}},\leavevmode\nobreak\ M\rightarrow\frac{M}{B^{+}M}.$ (9)
@@
The categories ${}^{\phantom{B}\mathllap{A}}_{\phantom{B}\mathllap{B}}\mathsf{mod}_{0}$ and ${{}^{\phantom{H}\mathllap{H}}_{\phantom{H}\mathllap{}}\mathsf{mod}}$ are monoidal, the monoidal structure on ${}^{\phantom{B}\mathllap{A}}_{\phantom{B}\mathllap{B}}\mathsf{mod}_{0}$ is the usual tensor product of comodules and $B$-bimodules while the monoidal structure on ${{}^{\phantom{H}\mathllap{H}}_{\phantom{H}\mathllap{}}\mathsf{mod}}$ is given by the usual tensor product of comodules.
@@ 2.3. Differential calculi and metrics
If $(g,(\leavevmode\nobreak\ ,\leavevmode\nobreak\ ))$ is a metric on $B$, then by [BM4, Lemma 1.16], $g$ is central, i.e, $bg=gb$ for all $b\in B$. This implies that the map $\operatorname{coev}_{g}$ is actually $B$-bilinear. The following well-known characterization (see page 311 of [BM4]) of metrics will be used throughout the article repeatedly.
@@
If $(\Omega^{\bullet},\wedge,d)$ is a *-differential calculus on *-algebra $B$ and $(g,(\leavevmode\nobreak\ ,\leavevmode\nobreak\ ))$ is metric on $\Omega^{1}$ then $g^{\dagger}=g$ if and only if $(\omega,\eta)=(\eta^{*},\omega^{*})^{*}$ for all $\omega,\eta\in\Omega^{1}$.
@@ 2.4. Complex structures
Following the work of Polishchuk and Schwarz in [PS], the notion of complex structures for noncommutative differential calculus have been studied by many mathematicians, see for example, [KLvS], [BS], and references therein. We will work with the following definition introduced in [OB1] where $\mathbb{N}_{0}$ will denote the set $\mathbb{N}\cup\{0\}.$
@@ 2.5. Bar Categories
A bar category is a monoidal category $(\mathcal{C},\otimes,1_{\mathcal{C}})$ together with a functor $\mathrm{bar}:\mathcal{C}\to\mathcal{C}$ (written as $X\mapsto\overline{X}$), a natural equivalence $\mathrm{bb}:\mathrm{id}_{\mathcal{C}}\to\mathrm{bar}\circ\mathrm{bar}$ between the identity and the bar $\circ$ bar functors on $\mathcal{C}$, an invertible morphism $\star:1_{\mathcal{C}}\to\overline{1_{\mathcal{C}}}$ and a natural equivalence $\Upsilon$ between $\mathrm{bar}\circ\otimes$ and $\otimes\circ(\mathrm{bar}\times\mathrm{bar})\circ\mathrm{flip}$ from $\mathcal{C}\times\mathcal{C}$ to $\mathcal{C}$
@@
and moreover, the following equations hold:
@@
An object $X$ in a bar category is called a star object if there is a morphism $\star_{X}:X\to\overline{X}$ such that $\overline{\star_{X}}\circ\star_{X}=\mathrm{bb}_{X}$.
@@ 2.6. Connections
([BM4]) Suppose $(g,(\leavevmode\nobreak\ ,\leavevmode\nobreak\ ))$ is a metric on the space of one-forms $\Omega^{1}$ of a differential calculus on an algebra $B$ in the sense of Definition 2.9. Then a bimodule connection $(\nabla,\sigma)$ is said to be compatible with $(g,(\leavevmode\nobreak\ ,\leavevmode\nobreak\ ))$ if $\nabla g=0$, where, by an abuse of notation, we have denoted the connection on $\Omega^{1}\otimes_{B}\Omega^{1}$ defined by (16) by $\nabla$ again.
@@ 3. Construction of metrics on quantum homogeneous spaces
In this section, we start by classifying covariant metrics on a quantum homogeneous space of the form $B=A^{\mathrm{co}(H)},$ where $A$ is a Hopf $\ast$-algebra and $H$ is a compact quantum group algebra (see Subsection 8.1). This classification is proved in Theorem 3.2. We will use this result to classify covariant metrics on the space of one-forms of the Heckenberger–Kolb calculus in Theorem 6.12. In the second subsection, we work in the set up of covariant complex structures on a quantum homogeneous space as above. Here, we construct a family of covariant metrics $(g,(\leavevmode\nobreak\ ,\leavevmode\nobreak\ ))$ which are real in the sense of Definition 2.12 and satisfies the condition $(\omega,\eta)=0$ if $\omega,\eta\in\Omega^{(1,0)}$ or if $\omega,\eta\in\Omega^{(0,1)}$.
@@ 3.1. The moduli space of metrics
So let us assume that $H$ is a compact quantum group algebra and moreover, $\Omega^{1}$ be the space of one-forms of a covariant $\ast$-differential calculus on $B$ such that $\Omega^{1}$ is an object of the category ${}^{\phantom{B}\mathllap{A}}_{\phantom{B}\mathllap{B}}\mathsf{mod}_{0}$. We will use the notation $V$ to denote $\Phi(\Omega^{1})$. Since the map $\star_{\Omega^{1}}$ defined in (14) is an isomorphism in ${}^{\phantom{B}\mathllap{A}}_{\phantom{B}\mathllap{B}}\mathsf{mod}_{0}$, we know that $\Phi(\star_{\Omega^{1}})$ induces an isomorphism in ${{}^{\phantom{H}\mathllap{H}}_{\phantom{H}\mathllap{}}\mathsf{mod}}$ from $V$ to $\overline{V}$ via Lemma 8.7. Moreover, if we fix an $H$-invariant inner product on $V$ (conjugate-linear on the right) in the sense of (58), then we have the isomorphism $\psi:\overline{V}\to\operatorname{Hom}(V,\mathbb{C})$ in ${{}^{\phantom{H}\mathllap{H}}_{\phantom{H}\mathllap{}}\mathsf{mod}}$ from Lemma 8.1. Hence, we have a right dual pairing of $V$ with itself defined by the following morphisms in ${{}^{\phantom{H}\mathllap{H}}_{\phantom{H}\mathllap{}}\mathsf{mod}}$:
@@
Now, by Takeuchi’s equivalence (Theorem 2.6), we have morphisms $\operatorname{ev}:\Omega^{1}\otimes_{B}\Omega^{1}\to B$ and $\operatorname{coev}:B\to\Omega^{1}\otimes_{B}\Omega^{1}$ in ${}^{\phantom{B}\mathllap{A}}_{\phantom{B}\mathllap{B}}\mathsf{mod}_{0}$ where $\Phi(\operatorname{ev})$ and $\Phi(\operatorname{coev})$ are defined by (17) and (18) respectively and moreover, the pair $(\operatorname{ev},\operatorname{coev})$ satisfies (2). Thus, by Remark 2.10, we immediately have the following result.
@@
Since $H$ is a compact quantum group algebra, the set of all $A$-covariant metrics on $\Omega^{1}$ is non-empty by Proposition 3.1. Consider the monoidal category ${}^{\phantom{B}\mathllap{A}}_{\phantom{B}\mathllap{B}}\mathsf{mod}_{0}$. Then by a combination of Remark 2.10 and Lemma 2.1, it follows that the space of $A$-covariant metrics is acted upon freely and transitively by the automorphism group of $\Omega^{1}$ in the category ${}^{\phantom{B}\mathllap{A}}_{\phantom{B}\mathllap{B}}\mathsf{mod}_{0}$. However, by Takeuchi’s equivalence, this group is isomorphic to the automorphism group of $\Phi(\Omega^{1})$ in the category ${{}^{\phantom{H}\mathllap{H}}_{\phantom{H}\mathllap{}}\mathsf{mod}}$.
@@ 3.2. Metrics on quantum homogeneous spaces with a complex structure
Using the fact that $MB^{+}=B^{+}M$ for any object $M$ in ${}^{\phantom{B}\mathllap{A}}_{\phantom{B}\mathllap{B}}\mathsf{mod}_{0}$ (see Definition 2.5), it can be easily checked that the map $\operatorname{ev}_{V^{(1,0)}}$ is well-defined. Moreover, the definition of $\operatorname{coev}_{V^{(1,0)}}$ is independent of the choice of the orthonormal basis.
@@
as $(\leavevmode\nobreak\ ,\leavevmode\nobreak\ )_{\Omega^{1}}=(\leavevmode\nobreak\ ,\leavevmode\nobreak\ )^{1,0}\oplus(-\lambda^{-1})(\leavevmode\nobreak\ ,\leavevmode\nobreak\ )^{0,1}\oplus 0\oplus 0$ and $\operatorname{coev}_{\Omega^{1}}=\operatorname{coev}^{1,0}\oplus(-\lambda)\operatorname{coev}^{0,1}\oplus 0\oplus 0$, where $(\leavevmode\nobreak\ ,\leavevmode\nobreak\ )^{1,0}$, $\operatorname{coev}^{1,0}$ and $(\leavevmode\nobreak\ ,\leavevmode\nobreak\ )^{0,1},\operatorname{coev}^{0,1}$ are as defined in Corollary 3.4 and Remark 3.5 respectively. If we denote $\operatorname{coev}_{\Omega^{1}}(1)$ by the symbol $g$, then $(g,(\leavevmode\nobreak\ ,\leavevmode\nobreak\ )_{\Omega^{1}})$ defines an $A$-covariant real metric on $\Omega^{1}$ (see Definition 2.12) which satisfies the condition
@@ 4.1. Construction of Hermitian metrics on quantum homogeneous spaces
Since the element $\sum^{n}_{i=1}e_{i}\otimes_{B}e^{i}$ is the inverse image of $\mathrm{id}_{\Omega^{1}}$ under the canonical isomorphism $\Omega^{1}\otimes_{B}\prescript{}{B}{\operatorname{Hom}(\Omega^{1},B)}\rightarrow\prescript{}{B}{\operatorname{Hom}(\Omega^{1},\Omega^{1})},$ the definition of $g$ is independent of the choice of the dual basis. It can be easily checked that $(g,(\leavevmode\nobreak\ ,\leavevmode\nobreak\ ))$ is a covariant metric on $\Omega^{1}.$
@@
Suppose $(\Omega^{(\bullet,\bullet)},\wedge,\partial,\overline{\partial})$ is an $A$-covariant complex structure on a quantum homogeneous space $B=A^{\mathrm{co}(H)}$, where $A$ is a Hopf $\ast$-algebra and $\Omega^{1}$ is an object of ${}^{\phantom{B}\mathllap{A}}_{\phantom{B}\mathllap{B}}\mathsf{mod}_{0}$. Moreover, we assume that $\Omega^{(1,0)}$ and $\Omega^{(0,1)}$ are non isomorphic simple objects in ${}^{\phantom{B}\mathllap{A}}_{\phantom{B}\mathllap{B}}\mathsf{mod}_{0}$.